OpenStudy (anonymous):

Khalid drove 66 miles on 3 gallons of gasoline. Which is the number of miles he could drive on 12 gallons of gasoline? a) 36 b) 198 c) 264 d) 792

OpenStudy (lgbasallote):

use ratio and proportion \[\frac{66}{3} = \frac{x}{12}\] solve for x

OpenStudy (anonymous):

hint 12=3*4